Gypsum is hydrated calcium sulphate (CaSO4.xH2O). 5.65 g of anhydrous calcium sulphate is obtained by heating 7.15 g of gypsum. Determine the empirical formula of gypsum.

solve this


1
Expert's answer
2021-03-24T06:49:48-0400

write out the equation first,assume that the number of water is x

CaSO4.xH2O(136+18x)=CaSO4(136)+xH2O(18x)


The difference between anhydrous calcium sulphate and x water is

7.15- 5.65 = 1.5g

Therefore the mass of water is 1.5g

136/5.65=18x/5.0


5.65g divided by 136g/mol equals to 0.042mol

7.15g divided by 0.042mol equals to 172.1 g/mol

172.1g/mol minus 136g/mol equals to 36.1 g/mol

36.1g/mol divided by 18g/mol equals to 2

Therefore empirical formula of gypsum is CaSO4·2H2O